OpenStudy (anonymous):

If a plane can travel 460 miles per hour with the wind and 400 miles per hour against the wind, find the speed of the plane without a wind and the speed of the wind. The speed of the plane in still air is __ miles per hour. The speed of the wind is __ miles per hour.

OpenStudy (anonymous):

for the speed of the plane simply (460+400)/2 and for the speed of the wind (460 -400)/2

OpenStudy (anonymous):

see actually here u have to setup equations let the actual speed of the plane be x miles/h and that of the wind be y miles/h with the wind combined speed becomes (x+y)miles/h and against the wind it is (x-y)miles/h as for here it is given that x+y=460 and x-y =400 hence u just need to solve these equations

OpenStudy (anonymous):

is it $460

OpenStudy (anonymous):

?

OpenStudy (anonymous):

speed of plane (460+400)/2 miles/hr = 430 miles /hr

OpenStudy (anonymous):

and that of wind = 460-430= 30 miles/hr

OpenStudy (anonymous):

oh i added them (slaps fore head) thxs

OpenStudy (anonymous):

yw